### v1

# Machine Learning Symmetry Discovery for Classical Mechanics

Abstract
In this study, we propose a data-driven, deep-learning-based Machine-Learning Symmetry Discovery (MLSD) algorithm to automate the discovery of continuous Lie group symmetries in classical mechanical systems from their time-evolution trajectory data. MLSD uses neural networks (NNs) to predict conserved physical quantities that implement symmetry transformations of the phase space coordinates. After training, MLSD is able to identify the Lie algebra, particularly non-abelian ones, as indicated by the Lie algebra structure coefficients. To demonstrate the effectiveness of the MLSD method, we applied it to simulated data from the classical three-dimensional Kepler problem and the harmonic oscillator. The results show that the algorithm successfully identified the hidden symmetry groups of both systems.
I Introduction
The term symmetry originates from the Greek words syn (meaning “same”) and metron (meaning “measure”). At its core, symmetry reflects a fundamental concept in our understanding of the natural world: the ability to preserve identical measurable properties under transformations or dynamics. Much of the progress in physics as a science has come from uncovering symmetries within physical systems, deepening our insight into the laws governing the universe[1, 2, 3, 4].
While the discovery of symmetries is invaluable, it is often a challenging task that typically requires the sophisticated expertise of physicists. To address this challenge, advancing field of machine learning[5, 6, 7] offers a promising solution for facilitating the comprehensive discovery of symmetries in complex systems. One significant advantage of machine learning in this context is its ability to learn and compress from large volumes of data. This data-driven approach can sometimes surpass human intuition, as machine learning algorithms can identify patterns and correlations that may not be manifest to human researchers. By harnessing the vast amounts of data generated in scientific research, machine learning holds the potential to accelerate discovery and provide new insights into the fundamental laws and symmetries governing physical systems[8, 9, 10].
Previous work in this area has achieved significant success in uncovering continuous symmetries in classical mechanical systems[11, 12, 13, 14, 15, 16, 17]. However, a more challenging extension is to develop a systematic framework capable of automatically discovering possible continuous Lie group symmetries, particularly non-Abelian ones, from a dataset, rather than relying on an analytical form of the Hamiltonian. In this study, we propose a Machine-Learning Symmetry Discovery (MLSD) algorithm designed to explore all possible continuous Lie group symmetry transformations in classical mechanical systems using their time-evolution datasets. The MLSD algorithm identifies conservation laws that correspond to continuous symmetry transformations. After training, MLSD outputs a three-way tensor $f$, representing the Lie algebra structure coefficients of the discovered set of symmetry transformations. The generality of these structure coefficients enables MLSD to identify both Abelian and non-Abelian symmetries.
The paper is organized as follows: we first review the definition of continuous symmetries in classical systems, then explain each loss term designed in MLSD to discover symmetries using neural networks, and finally examine our proposed algorithm on two tasks: discovering $\mathrm{SO}(4)$ symmetry in the three-dimensional Kepler problem; and discovering $\mathrm{SU}(3)$ symmetry in the three-dimensional harmonic oscillator.
Figure 1: The Machine Learning Symmetry Discovery (MLSD) framework takes time evolution data from a classical system and feeds the canonical coordinates into neural networks to predict a physical quantity. By optimizing the loss function, the predicted observables converge to conserved quantities, and the corresponding structure coefficients reveal the underlying symmetry group.
II Methodology
Our task of symmetry discovery differs from those explored in previous studies. For example, [11, 14] focuses on identifying transformations that preserve time evolution trajectories, while [12, 13] examines whether specific proposed transformations can be learned by neural networks. In contrast, our approach seeks to uncover symmetry transformations that extend beyond preserving the shape of trajectories and instead focus solely on preserving energy, as similarly discussed in [15, 16, 17], without relying on prior human knowledge. In this section, we begin by reviewing the concept of symmetry in classical systems as it pertains to our study. We then introduce the architecture of the Machine Learning Symmetry Discovery (MLSD) framework, which is designed to identify these symmetries.
II.1 Continuous symmetry and symmetry group of a physical system
The time evolution dynamics of a physical system is encoded in the Hamiltonian $H({\bm{x}})$, which depends on the canonical coordinates ${\bm{x}}:=({\bm{q}},{\bm{p}})$, with ${\bm{q}}$ and ${\bm{p}}$ being the conjugate position and momentum variables. A continuous symmetry of the system is defined as the conservation of energy under continuous symmetry transformations generated by a conserved quantity. This can be expressed through the vanishing Poisson bracket commutator between $H({\bm{x}})$ and the conserved quantity $G_{i}({\bm{x}})$, where $i\in\{1,2,...,n\}$ and $n$ denote the dimension of the symmetry group. The identification of the continuous symmetry group (Lie group) follows directly from the Lie algebra structure coefficients $f^{ijk}$, which are obtained from the commutators between the quantities.
$\{H,G_{i}\}=0,\forall\ i\in\{1,2,...,n\}.$ (1)
$\{G_{i},G_{j}\}=\sum_{k}f^{ijk}G_{k},\forall\ i,j\in\{1,2,...,n\}.$ (2)
Here the Poisson bracket $\{A,B\}$ between two canonical functions $A({\bm{x}})$ and $B({\bm{x}})$ is defined by the following vector-matrix-vector multiplication
$\{A,B\}:=(\nabla_{\bm{x}}A)^{\intercal}\mathbf{J}(\nabla_{\bm{x}}B),$ (3)
where $\nabla_{\bm{x}}$ is the gradient operator, and the matrix
$\mathbf{J}=\begin{pmatrix}&\mathds{1}_{d\times d}\\ -\mathds{1}_{d\times d}&\end{pmatrix}$ (4)
characterizes the symplectic metric of the canonical coordinate ${\bm{x}}$ in the phase space.
Since the three-way tensor $f^{ijk}$ is anti-symmetric by definition, we can assign $G_{0}({\bm{x}}):=H({\bm{x}})$ together with $f^{0jk}=f^{i0k}=f^{ij0}=0$. Therefore Eq. (1) and Eq. (2) can be combined for compactness as
$\{G_{i},G_{j}\}=\sum_{k}f^{ijk}G_{k},\forall\ i,j\in\{0,1,2,...,n\}.$ (5)
The goal of discovering all possible continuous symmetries is equivalent to finding all independent quantity functions $G_{i}({\bm{x}}),\forall i\in\{0,1,2,...,n\}$, along with the structure coefficient $f^{ijk}$ that define the symmetry group. The special $G_{0}$ component corresponds to the Hamiltonian $H$, relating to energy conservation. Its dependence $H({\bm{x}})$ on ${\bm{x}}$ can be learned from a series of time evolution data ${\bm{x}}(t)$, assuming the underlying principle of Hamiltonian dynamics of classical mechanics.
II.2 Learning continuous symmetry transformations
Considering a classical system in spatial dimension $d$, let ${\bm{X}}=\{{\bm{x}}(t)\in\mathbb{R}^{2d}:\dot{{\bm{x}}}=\{{\bm{x}},H\}\}$ be a dataset of time series samples, where each sample is a trajectory of the canonical coordinate ${\bm{x}}(t)$ following the underlying Hamiltonian dynamics driven by $H$. The data can be collected from observation, and the Hamiltonian $H$ may not be known to us. The goal is to uncover the continuous Lie group symmetries of such a classical dynamical system as defined by Eq. (5). To achieve this, we propose the Machine-Learning Symmetry Discovery (MLSD) algorithm, which uses neural networks (NNs) to parametrize each of the conserved quantities in Eq. (5) as $G_{i,\theta}$, where $\theta$ represents the learnable parameters in neural networks.
The Noether’s theorem states that these conserved quantities $G_{i,\theta}$ are also symmetry generators of the corresponding continuous symmetry. An infinitesimal symmetry transformation of ${\bm{x}}$ is given by the gradient of $G_{i,\theta}$ with respect to ${\bm{x}}$:
$\begin{split}{\bm{x}}^{\prime}=&{\bm{x}}+\{{\bm{x}},G_{i,\theta}\}\epsilon+\mathcal{O}(\epsilon^{2})\\ =&{\bm{x}}+\mathbf{J}(\nabla_{{\bm{x}}}G_{i,\theta})\epsilon+\mathcal{O}(\epsilon^{2}).\end{split}$ (6)
where $\epsilon$ denotes an infinitesimal variation and $\mathbf{J}$ is the symplectic metric defined in Eq. (4). Based on this principle, we can parametrize a conserved quantity $G_{i,\theta}$ and use its gradient, $\nabla_{{\bm{x}}}G_{i,\theta}$, to generate a symmetry transformation $\delta{\bm{x}}=\mathbf{J}(\nabla_{\bm{x}}G_{i})\epsilon$. In contrast to prior studies that directly model the symmetry transformation $\delta{\bm{x}}$ by neural networks[11, 12, 13, 14, 16], our approach of modeling $\delta{\bm{x}}$ indirectly through gradients of scalar functions $G_{i,\theta}$ is simpler, while inherently respecting the curl-free constraint $\nabla_{{\bm{x}}}\times(\nabla_{{\bm{x}}}G_{i,\theta})=0$ that should otherwise be imposed on $\delta{\bm{x}}$ as well.
  • Hamiltonian learning
    Let us start from learning the Hamiltonian of the classical system. Using (6), we expect $G_{0,\theta}$ can generate one step time evolution by $\{{\bm{x}},G_{0,\theta}\}=\dot{{\bm{x}}}$. Hence, the Hamiltonian of the sytem can be learned by minimizing the mean-square loss function:
    $\mathcal{L}_{\text{H}}=\mathop{\mathbb{E}}_{{\bm{x}}(t)\in{\bm{X}}}\int\big{\|}\mathbf{J}\nabla_{{\bm{x}}}G_{0,\theta}-\dot{{\bm{x}}}\big{\|}^{2}\mathrm{d}t.$ (7)
    where the trajectory ${\bm{x}}(t)$ is sampled over the time series dataset ${\bm{X}}$. In practice, the time derivative is approximated by $\dot{{\bm{x}}}(t)\simeq\frac{{\bm{x}}(t+\epsilon)-{\bm{x}}(t-\epsilon)}{2\epsilon}$ along the trajectory for some small $\epsilon$. It should be understood that the integrand is time dependent along the trajectory ${\bm{x}}(t)$, and the time $t$ integrates over the trajectory sampled from the dataset ${\bm{X}}$.
  • Symmetry discovery
    The symmetry discovery process begins by inputting a conjectured dimension $n$ of the Lie group. Therefore, $n+1$ individual neural networks are used to predict each conserved quantity. The structure coefficient $f^{ijk}$ is parametrized to be anti-symmetric, defined as $f^{ijk}=\theta^{ijk}-\theta^{ikj}+\theta^{jki}-\theta^{jik}+\theta^{kij}-\theta^{kji}$, for all $i,j,k\neq 0,$ with the parameter $\theta$ being optimizable. In order to recognize the symmetry group formed by all $G_{i,\theta}$, we train on the dataset ${\bm{X}}$ with the following mean square loss:
    $\begin{split}\mathcal{L}_{\text{G}}=\mathop{\mathbb{E}}_{{\bm{x}}(t)\in{\bm{X}}}\frac{1}{(n+1)^{2}}\sum_{i,j=0}^{n}\int\Big{\|}&(\nabla_{{\bm{x}}}G_{i,\theta})^{\intercal}\mathbf{J}(\nabla_{{\bm{x}}}G_{j,\theta})\\ &-\sum_{k=0}^{n}f^{ijk}G_{k,\theta}\Big{\|}^{2}\mathrm{d}t.\end{split}$ (8)
    It should be understood that the integrand is evaluated at instantaneously at each time $t$ and then integrated over the trajectory ${\bm{x}}(t)$ sampled from the dataset ${\bm{X}}$.
In addition to $\mathcal{L}_{\text{H}}$ and $\mathcal{L}_{\text{G}}$, we introduce a third term into the loss function to maximize the linear independence of the set of vectors $\nabla_{{\bm{x}}}G_{i,\theta}:i\in\{1,2,...,n\}$. Otherwise, the algorithm might converge to trival solutions where several $G_{i,\theta}$ correspond to equivalent conserved quantities, which we should try to avoid. For this purpose, we introduce the following $n\times 2d$ matrix
$\mathbf{M}({\bm{x}})=(\mathbf{J}(\widehat{\nabla_{{\bm{x}}}G_{0,\theta}}),\widehat{\nabla_{{\bm{x}}}G_{1,\theta}},...,\widehat{\nabla_{{\bm{x}}}G_{n,\theta}})^{\intercal},$ (9)
where each $\widehat{\nabla_{{\bm{x}}}G_{i,\theta}}=\nabla_{{\bm{x}}}G_{i,\theta}/{\lVert\nabla_{{\bm{x}}}G_{i,\theta}\rVert}$ denotes the normalized column vector. The linear independence of the learned transformations in Eq. (6) is equivalent to non-vanishing product of the top $\min(2d,n)$ [^1] If $12d\leq n+$ 1, then finite value of $\mathcal{L}_{\text{I}}$ indicates all the vectors in $\mathbf{M}({\bm{x}})$ have fully spanned the $2d$ dimension space; otherwise, the $n+1$ vectors in $\mathbf{M}({\bm{x}})$ are linearly independent to each other. eigenvalues of $\mathbf{M}({\bm{x}})\mathbf{M}^{\intercal}({\bm{x}})$. Denoting the eigenvalues of matrix $\mathbf{M}({\bm{x}})\mathbf{M}^{\intercal}({\bm{x}})$ as $\lambda^{i}({\bm{x}})$, the third term is written as:
$\mathcal{L}_{\text{I}}=-\mathop{\mathbb{E}}_{{\bm{x}}(t)\in{\bm{X}}}\int\frac{1}{N}\text{log}\Big{(}\prod_{i=1}^{\min(2d,n)}\lambda^{i}({\bm{x}})\Big{)}\mathrm{d}t.$ (10)
Since all the vectors defined in matrix $\mathbf{M}({\bm{x}})$ have been normalized, the condition $\prod_{i=1}^{\min(2d,n)}\lambda^{i}({\bm{x}})>0$ indicates maximal independence of the learned symmetry transformations. Consequently, in practice, the independence term in Eq. (10) is scaled by a very small factor when added to the overall loss function to prevent blowing up the total loss function.
In summary, the overall loss function is given by:
$\mathcal{L}=\mathcal{L}_{\text{H}}+\alpha\mathcal{L}_{\text{G}}+\beta\mathcal{L}_{\text{I}}.$ (11)
with the factor $\beta\ll\alpha\simeq 1$.
After training, if the given dataset ${\bm{X}}$ corresponds to a system with continuous $n$ dimensional symmetry, the first two loss terms, Eq. (7) and Eq. (8), should converge to zero, while the third term, Eq. (10), remains finite. Conversely, if the system lacks an $n$ dimensional symmetry, the first two loss terms will remain finite regardless of how small $\beta$ is set. The specific type of symmetry group can be identified by extracting the structure coefficients $f^{ijk}$ , as detailed in Appendix B.
Hamiltonian Conserved quantities Lie Algebra Lie Group
[…]
Table 1: Summary of hidden symmetries in harmonic oscillator and Kepler problem.
III Empirical results
In this section, we aim to implement MLSD to discover symmetries in the harmonic oscillator and Kepler problem, summarized in Tab. 1. A detailed analysis of $\mathrm{SU}(3)$ symmetry of harmonic oscillator and $\mathrm{SO}(4)$ symmetry of Kepler problem is provided in the Appendix A. We demonstrate that MLSD can successfully identify the hidden symmetries in both systems. Additionally, in the harmonic oscillator task, we use quadratic parametrization to show that MLSD can explicitly reconstruct the 8 Gell-Mann matrices that form the $\mathfrak{su}(3)$ Lie algebra. Code is available at [19].
III.1 Kepler problem
Following the pipeline outlined in Sec. II, we tested multiple conjectured symmetry group dimensions, $n=6$, $n=10$, and $n=15$  [^2] An $\mathrm{SO}(3)$ symmetry with $n=3$ is obvious since the Kepler Hamiltonian preserves the angular momentum with rotation symmetry., corresponding to $\mathrm{SO}(4)$, $\mathrm{SO}(5)$, and $\mathrm{SO}(6)$, respectively. We then trained the neural networks on the simulation dataset using the loss function in (11) with $\alpha=1$ and $\beta=1\times 10^{-4},2\times 10^{-4}...,10^{-3}$ as shown in Fig. 3.
Figure 2: Examples of time evolution driven by the learned Hamiltonian. Each sub-figure compares the time evolution generated by the true Hamiltonian $H_{\text{Kep}}$ (dashed gray) and the learned Hamiltonian $G_{0,\theta}$ (blue), both starting from the same initial condition ${\bm{x}}(t=0)$. The evolution trajectories are plotted after projecting the $3d$ variables onto a $2d$ plane.
A straight forward method to evaluate the Hamiltonian learning is by reproducing the time evolution trajectories, as illustrated in Fig. 2. Starting from the same initial condition ${\bm{x}}(t=0)$, the well-trained $G_{0,\theta}$ is able to reproduce the time evolution Eq. (6) highly identical to that of $H$ for a finite amount of time, before losing track due to potentially accumulating errors.
By tracking the converged values of each term in Eq. (11) for $n=6$, $n=10$, and $n=15$, we find that only the $n=6$ conjecture with $\beta=1\times 10^{-4},2\times 10^{-4},3\times 10^{-4}$ can minimize $\mathcal{L}_{\text{G}}$ to zero while keeping $\mathcal{L}_{\text{I}}$ finite, as shown in Fig. 3. This suggests a potential symmetry group of dimension $n=6$ that is worth further investigation.
Figure 3: The sub-figures (a), (b), and (c) show $\mathcal{L}_{\text{G}}$ versus training epochs for $n=6$, $n=10$, and $n=15$. (d) shows the converged $\mathcal{L}_{\text{G}}$ versus $\beta$. The $\mathcal{L}_{\text{G}}(n=6)$ with $\beta=1\times 10^{-4},2\times 10^{-4},3\times 10^{-4}$ are able to converge to near zero comparing to other trials.
To further identify the learned symmetry group, we diagonalize the Killing form matrix $\mathbf{B}=\mathbf{U}\mathbf{D}\mathbf{U}^{T}$, whose entries defined as $B^{il}=\sum_{k,j}f^{ijk}f^{lkj}$, where $f^{ijk}$ is the machine-learned structure coefficient in Eq. (2). Example results yields $\mathbf{D}=\text{diag}(-3.40,-3.44,-3.47)\oplus\text{diag}(-1.22,-1.20,-1.17)$. This pattern is consistent with theoretical expectations of the decoupling of the two $\mathfrak{su}(2)$ algebras, as $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$. Further projecting the structure coefficient using $\mathbf{U}$ reproduces the $\mathfrak{so}(4)$ algebra as shown in Fig. 4. A detailed explanation of the method used to identify this symmetry group is provided in Appendix B.
Figure 4: The trained structure coefficient of Kepler problem.
III.2 Harmonic oscillator
We apply the same method to test multiple conjectured symmetry group dimensions $n=8$, $n=10$, $n=15$. The results of Hamiltonian learning are presented in Fig. 5. The learning task is notably easier in the harmonic oscillator system compared to the Kepler problem, as the Hamiltonian and conserved quantities in the harmonic oscillator have a simpler quadratic form in terms of ${\bm{p}}$ and ${\bm{q}}$.
At $n=8$, the 8 eigenvalues of the Killing form of the structure coefficients exhibit strong degeneracy, for example: $\mathbf{D}=\mathrm{diag}$($-18.48$, $-18.46$, $-18.44$, $-18.43$, $-18.43$, $-18.41$, $-18.40$, $-18.38$). The structure coefficient reconstructs the $\mathfrak{su}(3)$ Lie algebra as shown in the Fig. 6. This strongly indicates a hidden $\mathrm{SU}(3)$ symmetry in the harmonic oscillator system.
Figure 5: Examples of time evolution driven by the learned Hamiltonian in harmonic oscillator system.
Figure 6: The trained structure coefficient of harmonic oscillator, obtained through both quadratic and NN parametrizations, yield similar results that are identical to the $\mathfrak{su}(3)$ Lie algebra.
Figure 7: The sub-figures (a), (b), and (c) show $\mathcal{L}_{\text{G}}$ versus training epochs for $n=8$, $n=10$, and $n=15$. (d) shows the converged $\mathcal{L}_{\text{G}}$ versus $\beta$. The $\mathcal{L}_{\text{G}}(n=8)$ with $\beta=1\times 10^{-4},2\times 10^{-4},3\times 10^{-4},,4\times 10^{-4}$ are able to converge to near zero comparing to other trials.
We also tried the method that parameterizes 9 variables in quadratic forms as $G_{i}={\bm{x}}^{T}M_{i}{\bm{x}}$, where $i\in(0,1,2,...,8)$ and each $M_{i}$ is a real matrix of 6 by 6 that can be optimized. Here $G_{0}$ represents the Hamiltonian and others represent the 8 conserved quantities.
After minimizing the loss function (11), the 8 conserved quantities explicitly reconstruct the Gell-Mann basis Fig. 8.
Figure 8: Coefficient vectors of 8 quadratic quantities, each element represents the coefficient in the order of: $q_{1}q_{1}$, $q_{1}q_{2}$, $q_{1}q_{3}$, $q_{2}q_{2}$, $q_{2}q_{3}$, $q_{3}q_{3}$, $p_{1}q_{1}$, $p_{1}q_{2}$, $p_{1}q_{3}$, $p_{1}p_{1}$, $p_{1}p_{2}$, $p_{1}p_{3}$, $p_{2}q_{1}$, $p_{2}q_{2}$, $p_{2}q_{3}$, $p_{2}p_{2}$, $p_{2}p_{3}$, $p_{3}q_{1}$, $p_{3}q_{2}$, $p_{3}q_{3}$, $p_{3}p_{3}$.
IV Summary and Discussion
  • Symmetry discovery without prior knowledge
    In this study, we propose a data-driven, deep learning-based Machine Learning Symmetry Discovery algorithm designed to explore and analyze continuous symmetries in classical mechanical systems. Previous works[12, 13, 14, 11, 15, 16, 21, 22] have demonstrated methods to verify conjectured continuous symmetries of a given Hamiltonian. To make further improvements, our pipeline requires only the conjectured dimension of the symmetry group. The machine learning algorithm can then discover the symmetry transformations and subsequently identify the symmetry group, whether Abelian or non-Abelian, by analyzing the Killing form of the structure coefficients. This study addresses a crucial topic: discovering hidden symmetries within data itself using machine learning. Many continuous symmetries in physical systems are often obscure and difficult to uncover, even when the explicit form of the Hamiltonian is known. By leveraging deep learning, we aim to surpass human capabilities in identifying these symmetries. Our approach enables the algorithm to learn directly from the data, uncovering potential underlying symmetries that might otherwise remain hidden through traditional analysis. This represents a significant advancement in the application of machine learning to the discovery of complex patterns in physical systems.
  • Extension to quantum and many-body system
    Extending the idea of symmetry discovery to quantum many-body systems is a highly valuable area of study. Understanding the hidden symmetries in these complex systems can provide deeper insights into the exotic properties of quantum states, and guide the classification of different states of matter based on their symmetries. However, directly computing symmetries in arbitrarily large quantum systems is infeasible due to the exponential growth of the Hilbert space. Therefore, developing a more efficient framework that combines renormalization group techniques[23, 24, 25] with symmetry discovery is an important avenue for future research. This approach could potentially reduce the computational complexity while still capturing the essential symmetries of the system, making it a promising direction for exploring the rich structures within quantum many-body physics.
  • Limitations
    Many symmetry transformations driven by quantitys have highly nonlinear dependence on the canonical coordinates. If the target symmetry transformation exceeds the representational capacity of linear feed-forward neural networks, then more advanced neural networks [7, 26, 27, 28, 29] are required for symmetry discovery. Otherwise, the learning process may be hindered. Developing neural networks specifically tailored for AI-driven scientific tasks is crucial, as these tasks often involve complex structures and patterns that demand greater representational power and precision compared to general-purpose neural networks.
    The MLSD algorithm relies on an independence loss term $\mathcal{L}_{\text{I}}$ to encourage the search for linearly independent symmetry transformations. This independence loss plays a crucial role in guiding the model to discover distinct symmetries. However, despite being weighted by a small parameter $\beta$, this term can still introduce instability into the gradient descent process, making the optimization unstable and potentially causing the task to fail. A better design for the symmetry searching task is still worth exploring.
Acknowledgements.
We acknowledge helpful discussion with Rose Yu and Ziming Liu. WD and YZY are supported by the National Science Foundation (NSF) Grant DMR-2238360. ML is supported by UCSD Student Success Center 2024 Undergraduate Summer Research Award.
Appendix A Symmetry Analysis
  • Kepler problem
    We begin by defining several fundamental quantities. Let the coordinates be ${\bm{q}}=\{q_{1},q_{2},q_{3}\}$ and the momenta be ${\bm{p}}=\{p_{1},p_{2},p_{3}\}$. The Hamiltonian is given by
    $H=\frac{{\bm{p}}^{2}}{2}-\frac{1}{|{\bm{q}}|},$
    the angular momentum vector is
    ${\bm{L}}={\bm{q}}\times{\bm{p}},$
    and the Laplace–Runge–Lenz (LRL) vector is
    ${\bm{A}}={\bm{p}}\times{\bm{L}}-\frac{{\bm{q}}}{|{\bm{q}}|}.$
    We utilize Poisson brackets to derive the commutation relations, defined as
    $\{A,B\}=\frac{\partial A}{\partial q^{i}}\frac{\partial B}{\partial p_{i}}-\frac{\partial A}{\partial p^{i}}\frac{\partial B}{\partial q_{i}}.$
    Calculate commutation relations between Hamiltonian:
    $\{H,L_{i}\}=0,\quad\{H,A_{i}\}=0,\quad i\in\{1,2,3\}.$
    This indicates that both ${\bm{L}}$ and ${\bm{A}}$ are conserved quantities.
    Calculating the commutation relationships within and between ${\bm{L}}$ and ${\bm{A}}$:
    $\{L_{i},L_{j}\}=\epsilon^{ijk}L_{k},\quad\{L_{i},A_{j}\}=\epsilon^{ijk}A_{k},\quad\{A_{i},A_{j}\}=-2H\epsilon^{ijk}L_{k},\quad ijk\in\{1,2,3\}.$
    We define the scaled Laplace–Runge–Lenz vector ${\bm{N}}\equiv\frac{{\bm{A}}}{\sqrt{-2H}}$, and observe that ${\bm{N}}$ and ${\bm{L}}$ generate the $\mathfrak{so}(4)$ algebra from the commutation relations above. Furthermore, we introduce the vectors
[…]
    and
[…]
    The commutation relations for ${\bm{\mathcal{L}}}$ and ${\bm{\mathcal{R}}}$ are given by:
    $\{\mathcal{L}_{i},\mathcal{L}_{j}\}$ $=\epsilon^{ijk}\mathcal{L}_{k},$
    $\{\mathcal{R}_{i},\mathcal{R}_{j}\}$ $=\epsilon^{ijk}\mathcal{R}_{k},$
    $\{\mathcal{L}_{i},\mathcal{R}_{j}\}$ $=0.$
    These relations demonstrate that ${\bm{\mathcal{L}}}$ and ${\bm{\mathcal{R}}}$ each satisfy the commutation relations of the $\mathfrak{su}(2)$. Consequently, we have two decoupled $\mathfrak{su}(2)$, where
    $\mathfrak{su}(2)\oplus\mathfrak{su}(2)\cong\mathfrak{so}(4).$
  • Harmonic oscillator
    Using the same definitions of coordinate and momentum in the Kepler problem, we define the classical annihilation and creation functions as
    ${\bm{a^{\dagger}}}=\{a^{*}_{1},a^{*}_{2},a^{*}_{3}\}\quad\text{and}\quad{\bm{a}}=\{a_{1},a_{2},a_{3}\},$
    where $a_{i}=q_{i}+\mathrm{i}p_{i},\;i\in\{1,2,3\}$.
    The Hamiltonian $H$ is given by
    $H=\frac{1}{2}({\bm{q}}^{2}+{\bm{p}}^{2})=\frac{1}{2}{\bm{a^{\dagger}}}{\bm{a}}=\frac{1}{2}{\bm{a^{\dagger}}}\mathds{1}{\bm{a}}.$
    Note that $\{{\bm{a^{\dagger}}}A{\bm{a}},{\bm{a^{\dagger}}}B{\bm{a}}\}=-2\mathrm{i}{\bm{a^{\dagger}}}[A,B]{\bm{a}}$. By constructing the generators as $G_{i}=\frac{1}{2}{\bm{a^{\dagger}}}M_{i}{\bm{a}}$, we automatically have
    $\{H,G_{i}\}=-\frac{\mathrm{i}}{2}{\bm{a^{\dagger}}}[\mathds{1},M_{i}]{\bm{a}}=0.$
    We also want the structure constants to be anti-symmetric, leading to
    $\{G_{i},G_{j}\}=\frac{\mathrm{i}}{2}{\bm{a^{\dagger}}}[M_{i},M_{j}]{\bm{a}}=f^{ijk}\frac{1}{2}{\bm{a^{\dagger}}}M_{k}{\bm{a}}=f^{ijk}G_{k}.$
    In this case, it is natural to choose $\{M_{i}\}$ to be the Gell-Mann matrices, as the Gell-Mann matrices together with the identity matrix form a complete basis of 3 by 3 Hermitian matrices.
    The structure constants $f^{ijk}$ turn out to be anti-symmetric, with the following non-zero values and their permutations:
    $f^{123}=2,\quad f^{147}=f^{246}=f^{257}=f^{345}=1,\quad f^{156}=f^{367}=-1,\quad f^{456}=f^{678}=\sqrt{3}.$
Appendix B Basis transformation of Lie algebra
  • Linear transformation of the structure coefficient
    Considering two sets of Lie algebra basis ${\bm{G}}=(G_{i},...,G_{n})$ and ${\bm{e}}=(e_{1},...,e_{n})$ are related by a linear transformation by $M$:
    ${\bm{G}}=M{\bm{e}}$
    $\{e_{a},e_{b}\}=\epsilon_{ab}^{c}e_{c},\quad\forall a,b$
    $\{G_{i},G_{j}\}=f_{ij}^{k}G_{k},\quad\forall i,j$
    then the structure coefficient is related by
    $f_{ij}^{k}M_{k}^{c}e_{c}=\{M_{i}^{a}e_{a},M_{j}^{b}e_{b}\}=M_{i}^{a}M_{j}^{b}\{e_{a},e_{b}\}=M_{i}^{a}M_{j}^{b}\epsilon_{ab}^{c}e_{c}$
    $f_{ij}^{k}=M_{i}^{a}M_{j}^{b}\epsilon_{ab}^{c}(M^{-1})_{c}^{k}$
    $(M^{-1})_{a}^{i}(M^{-1})_{b}^{j}f_{ij}^{k}M_{k}^{c}=\epsilon_{ab}^{c}$
    Therefore, if there exists such a matrix $M$ that relates $f$ and $\epsilon$ through the above equality, then ${\bm{G}}$ and ${\bm{e}}$ can be identified as representations of the same Lie algebra.
  • Identify Lie group from diagonalizing the Killing form.
    There are two methods to identify the Lie group after training with MLSD. Suppose $f$ is the structure coefficient obtained from MLSD and $\epsilon$ represents a reference coefficient. The straightforward approach is to use gradient descent to find an optimal transformation matrix $M$ by minimizing the objective function: $\sum_{a,b,c}\|(M^{-1})_{a}^{i}(M^{-1})_{b}^{j}f_{ij}^{k}M_{k}^{c}-\epsilon_{ab}^{c}\|$. If such a matrix $M$ exists, than the two groups are isomorphism to each other.
    The second approach is to diagonalize the Killing form matrix defined as
    $\mathbf{B}_{i}^{l}=f^{k}_{ij}f^{j}_{lk}$
    $\mathbf{D}=\mathbf{U}^{-1}\mathbf{B}\mathbf{U}$
    to obtain $\mathbf{D}$. The eigenvalues of the Killing form matrix $\mathbf{B}$ will exhibit degeneracy corresponding to the sub-algebra structure of ${\bm{G}}$. For instance, in the Kepler problem, the symmetry algebra is $\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2)$ Consequently, diagonalizing the Killing form matrix of the trained structure coefficients $f$ should yield two sets of degenerate eigenvalues: three identical eigenvalues corresponding to one $\mathfrak{su}(2)$ sub-algebra and another three identical eigenvalues corresponding to the other $\mathfrak{su}(2)$ sub-algebra. This eigenvalue pattern reflects the direct sum structure of the algebra. In the case of the harmonic oscillator problem, the symmetry algebra is $\mathfrak{su}(3)$ , which cannot be decomposed into a direct sum of sub-algebras. Therefore, the eigenvalues of the Killing form matrix should exhibit a single set of eight identical values, reflecting the irreducibility of the $\mathfrak{su}(3)$ algebra.